OpenStudy (anonymous):

The equation y = x - 3 represents a line perpendicular to y = -x + 3. Is This TRUE or FALSE?

OpenStudy (precal):

true look at your slopes y=x-3 m=1 y=-x+3 m=-1 they are opposite reciprocals